OpenStudy (anonymous):

312 mL of 1.50 x 10^-2 M NaOH is used to titrate 225 mL of a monoprotic acid of unknown concentration to reach the equivalence point. What is the molarity of the acid? 1.50 x 10^-2 M 4.68 x 10^-3 M 2.08 x 10^-2 M 1.05 x 10^-3 M

OpenStudy (anonymous):

ok so you probably know this equation: c=n/V -> n=c*V using this equation you should calculate amount of NaOH in moles then you know from the text that it is monoprotonic acid so ratio of NaOH: X = 1:1 now that we know that we know that amount of X is equal to amount of NaOH so now we just calculate its concentration by dividing moles with its volume.

OpenStudy (anonymous):

and the answer is: 2.08 x 10^-2 M

OpenStudy (anonymous):

he he, you should open new question but you can look one question under yours, its the same question :) you are obviously solving same test :)

OpenStudy (anonymous):

oh lol that is really weird. there is a lot of my classmates that come here

OpenStudy (anonymous):

:)